# TuneNSearch: a hybrid transfer learning and local search approach for solving vehicle routing problems

### v2

@@ 4.1.2 Decoder details
Following the approach proposed by Kool et al. (2019), we apply an MHA layer followed by a SHA layer for the decoder. First, the decoder takes the initial node embeddings $x_{i}^{(L)}$ (we omit the $(L)$ term for better readability) and sets the keys and values for all $H$ heads of the MHA, as indicated in Equations 9 and  10:
@@
$v=k_{i}=W^{V}=Kx_{i},\forall i\in\left\{1,2,...,g\right\}$ (10)
@@
where the results are clipped within $[-C,C]$. Furthermore, the compatibility of infeasible nodes is set to $\infty$- ∞ to guarantee that only feasible solutions are generated. Lastly, the probability $p_{i},\forall i\in\left\{1,2,…,g\right\}$ of each node is computed through a softmax function, as indicated in Equation 13.
@@ 4.2 Model training
To train our model, we used the REINFORCE algorithm (Williams, 1992), which is a fundamental policy gradient method used in reinforcement learning. In particular, we use the REINFORCE with shared baselines algorithm, following the approach of POMO (Kwon et al., 2020), where multiple trajectories are sampled with $N$ different starting nodes. To this end, the total rewards $R(\tau^{1},…,\tau^{N})$ are calculated for each solution sampled. For a batch with $B$ different instances, gradient ascent is used to maximize the total expected return $J$, as indicated in Equation 14:
@@
Algorithm 1 outlines the REINFORCE with shared baselines algorithm in more detail. First, the policy network is initialized with a random set of parameters $\theta$. Then, training begins by looping over $E$ epochs, each comprising $T$ steps. At each step, a batch of $B$ random training instances is sampled, and for each instance, $N$ starting nodes are selected. For all experiments, we set $N$ as $g-1$, consistent with prior work (Kwon et al., 2020; Li et al., 2024). From these nodes, trajectories are sampled through the model architecture explained in Section 4.1. Then, the shared baseline is calculated (Equation 15), which is used to compute the policy gradients $\nabla_{\theta}J(\theta)$ (Equation 14). Lastly, the set of parameters $\theta$ is updated through gradient ascent, scaled by a learning rate $\eta$.
@@
5:        Select $N$ starting nodes for each instance $\mathcal{G}_{i}$, $i\in\{1,...,B\}$
@@ 4.3 Pre-training and fine-tuning process
With the architecture described above, we pre-train a backbone model for 100 epochs, similarly to the framework proposed by Lin et al. (2024). However, rather than training on TSP data, we use MDVRP instances. This choice is motivated by the higher complexity of the MDVRP, which enables the model to learn richer node representations in comparison to the TSP. Besides, pre-training on the MDVRP also offers advantages over pre-training on the CVRP by exposing the model to a multi-depot structure. As a result, our model generalizes as well as a CVRP pre-trained model on single-depot variants, while achieving significantly better performance on multi-depot tasks.
Additionally, in contrast to Lin et al. (2024), during pre-training we incorporate all dynamic features $D_{t}$ relevant across all VRP variants considered. Rather than introducing problem-specific modules, this approach allows us to fine-tune the model for each VRP variant without modifying the neural network architecture. This avoids the loss of potentially useful knowledge when transitioning between variants. The only structural change occurs when adapting the model for the TSP, as it is the simplest routing problem we address. The TSP differs from other VRP variants since it involves only node coordinates, with no demands, time windows, or additional constraints. For TSP-specific fine-tuning, we modify the encoder by removing the fully connected layer responsible for embedding customer features, leaving only depot (characterized solely by coordinates) and edge features.
For the fine-tuning phase, we load the parameters from the pre-trained MDVRP model and train it for an additional 20 epochs on randomly generated instances, tailored to each specific VRP variant. We use the same hyper-parameters as in the pre-training phase. The dynamics of the pre-training and fine-tuning stages of TuneNSearch are illustrated in Fig. 2. We note that besides the main variants described in Section 3.3, TuneNSearch can be extended to any combination of VRP constraints. For example, it can handle combinations like the MDVRP with time windows, open routes and backhauls, allowing it to tackle more complex routing problems.
@@ 4.4 Local search algorithm
Despite the recent interest in using neural-based techniques to solve VRPs, there remains a limited integration between operations research and machine learning methods in the current literature. Bridging this gap offers an opportunity to combine the strengths of both fields, enabling the development of more powerful and efficient algorithms. To refine the solutions obtained by the neural-based model, we designed an efficient local search algorithm employed after inference, inspired by different existing methods.
This algorithm involves applying various operators within restricted neighborhoods of predefined size. Following Vidal (2022), we define the neighborhood size as 20, which allows for an efficient exploration of the granular neighborhood. A granular neighborhood is a restricted subset of the solution space that is defined based on proximity criteria. Rather than evaluating moves across the entire problem domain, the search is confined to carefully selected neighborhoods of limited size, where meaningful improvements are more likely to be found. Specifically, the chosen operators apply moves restricted to node pairs $(a,b)$, where $b$ is one of the 20 closest nodes to $a$. All moves are evaluated within different neighborhoods, and any improvement in the cost function is immediately applied. The search procedure terminates once no further improvements to the cost function are possible, which occurs after all applicable operators and moves have been applied. We adopt the same set of operators as Wouda et al. (2024), which were selected due to their ability to effectively explore the solution space while maintaining computational efficiency. Each operator targets different aspects of the solution structure, ensuring a well-balanced and diverse set of moves to improve the solution quality. The chosen operators include:
@@
    (X, M)-exchange: involves swapping $X$ customers from a route, starting at a designated node, with a segment of $M$ customers (where $0\leq M\leq X$) starting at a different node. Importantly, the two segments must not overlap, ensuring that the exchange modifies the routes without duplicating any customers.
@@
    RELOCATE*: identifies and executes the most optimal relocation where an individual customer is removed from one route and inserted into the best possible position of another route.
@@
    SWAP*: evaluates and executes the most beneficial exchange of two customers between two routes, positioning each customer in the optimal location within the other route. Unlike a traditional swap, this operator exchanges two customers from different routes by inserting them into any position on the opposite route, rather than directly swapping them in place.
In more detail, the local search algorithm (see Algorithm 2) begins with an initial solution $\tau$, which is the best solution obtained from the E-GAT model. The algorithm assumes a fixed number of iterations $I$, and solutions are represented as sequences of visited nodes. The best distance ($BD$) is initialized as the cost of $\tau$, and the best solution ($BS$) is set to $\tau$. Before the iterative process starts, the algorithm explores the neighborhood of $\tau$ using the $Search$ function. This function applies all the operators and moves explained above to identify potential improvements to the solution. If a better solution is found during this step, both $BD$ and $BS$ are updated accordingly. After that, in each iteration, to escape local minimum, an offspring solution ($OS$) is generated using the $𝑜𝑣𝑒𝑟Cross$ function. This function takes as parents the current $BS$ and a randomly generated solution, created with the $MakeRandom$ function. Here, we use the selective route exchange crossover operator proposed by Nagata and Kobayashi (2010). Then, the neighborhood of $OS$ is explored through the $Search$ function. If the resulting solution is infeasible, the algorithm repairs it using the $Fix$ function, which applies the search operators again while considering only feasible moves. Finally, if the new solution’s cost is better than the current $BD$, the algorithm updates $BD$ and $BS$. This process is repeated for $I$ iterations.
Fig. 3 depicts an overview of the complete inference process, coupled with the local search algorithm. The process begins with the trained E-GAT model solving a given testing instance, generating an initial solution represented as sub-tours (or routes). This solution is then improved by the local search algorithm, which explores the solution space. Finally, the best overall solution found is returned.
Figure 3: Overview of the inference process of TuneNSearch.
@@
1:  $BD\leftarrow Cost(\tau)$
2:  $BS\leftarrow\tau$
3:  $\tau^{\prime}\leftarrow Search(\tau)$

### v3

@@ 4.1.2 Decoder details
Following the approach proposed by Kool et al. (2019), we apply an MHA layer followed by a SHA layer for the decoder. First, the decoder takes the initial node embeddings $x_{i}^{(L)}$ (we omit the $(L)$ term for better readability) and sets the keys and values for all $H$ heads of the MHA, as indicated in Equations 8 and  9:
@@
$k_{i}=W^{K}x_{i},\forall i\in\left\{1,2,...,g\right\}$ (9)
@@
where the results are clipped within [C,C][-C,C]. Furthermore, the compatibility of infeasible nodes is set to -$\infty$ to guarantee that only feasible solutions are generated. Lastly, the probability $p_{i},\forall i\in\left\{1,2,…,g\right\}$ of each node is computed through a softmax function, as indicated in Equation 12.
@@
This process continues indefinitely, with the decoder selecting one node at a time, until a complete solution is constructed for each instance.
@@ 4.2 Model training
To train our model, we used the REINFORCE algorithm (Williams, 1992), which is a fundamental policy gradient method used in reinforcement learning. In particular, we use the REINFORCE with shared baselines algorithm, following the approach of POMO (Kwon et al., 2020), where multiple solutions/trajectories are sampled with $N$ different starting nodes. A trajectory refers to the full sequence of actions or decisions that the model takes to construct a complete solution. In the context of VRPs, a trajectory can essentially be interpreted as a solution. This means that, in a batch with $B$ different instances, the model simultaneously constructs $N$ trajectories per instance, generating $B\times N$ solutions in parallel. For each trajectory, the first action is deterministically chosen, after which the decoder will continue to build each trajectory as explained before, according to the calculated node probabilities. The rationale behind this design is to encourage diversity in the generated solutions, since each customer has the opportunity to serve as a starting point. As a result, the model explores a broader range of possible solutions, improving its coverage of the solution space.
Once all $B\times N$ trajectories have been constructed, we compute the total rewards $R(\tau^{1},…,\tau^{N})$ for each solution, and use gradient ascent to maximize the total expected return $J$, as indicated in Equation 13:
@@
Algorithm 1 outlines the REINFORCE with shared baselines algorithm in more detail. First, the policy network is initialized with a random set of parameters $\theta$. Then, training begins by looping over $E$ epochs, each comprising $T$ steps. At each step, a batch of $B$ random training instances is sampled, and for each instance, $N$ starting nodes are selected. For all experiments, we set $N$ as g1g-1, consistent with prior work (Kwon et al., 2020; Li et al., 2024). From these nodes, trajectories are sampled through the model architecture explained in Section 4.1. In our implementation, node indices are assigned such that depots occupy the first positions, followed by customers (e.g., 0, 1 and 2 for a problem with three depots). Like prior work (Li et al., 2024), depot 0 is treated as the implicit starting point of the first route unless the deterministic starting node itself corresponds to another depot (e.g., index 1 or 2). For example, in a problem with 3 depots and 50 customers, we set $N=52$: two trajectories start from depots with indices 1 and 2, while the remaining 50 trajectories start from customers with indices 3–52, all initially departing from depot 0. Although one could consider all depot–customer pairs as starting states (yielding 150 trajectories in this case), the computational cost would be prohibitive. Importantly, initializing multiple trajectories at depot 0 does not bias the solutions towards a single depot. The reason is that depot assignment is not fixed by the initialization: throughout decoding, the model explicitly decides when to close a route and from which depot the next vehicle should depart. Thus, even if the first route in a trajectory begins at depot 0, subsequent routes may start at other depots whenever the policy deems it as beneficial.
After this, the shared baseline is calculated (Equation 14), which is used to compute the policy gradients $\nabla_{\theta}J(\theta)$ (Equation 13). Lastly, the set of parameters $\theta$ is updated through gradient ascent, scaled by a learning rate $\eta$.
@@
5:   Select NN starting nodes for each instance $\mathcal{G}_{i}$, $i\in\{1,...,B\}$
@@ 4.3 Pre-training and fine-tuning process
With the architecture described above, we pre-train a backbone model for 100 epochs, similarly to the framework proposed by Lin et al. (2024). However, rather than training on TSP data, we use MDVRP instances. This choice is motivated by the higher complexity of the MDVRP, which enables the model to learn richer node representations in comparison to the TSP. Pre-training on MDVRP data also offers advantages over using CVRP data, since the model is exposed to multi-depot structures.
Additionally, in contrast to Lin et al. (2024), during pre-training we incorporate all dynamic features $D_{t}$ relevant across all VRP variants considered. Rather than introducing problem-specific modules, this approach allows us to fine-tune the model for each VRP variant without modifying the neural network architecture. This avoids the loss of potentially useful knowledge when transitioning between variants. The only structural change occurs when adapting the model for the TSP, as it is the simplest routing problem we address. The TSP differs from other VRP variants since it involves only node coordinates, with no demands, time windows, or additional constraints. In fact, the only necessary features to encode a TSP instance are the 2D coordinates of each node in the problem and the edge distances between each pair of nodes. Thus, for TSP-specific fine-tuning, we make the following adjustments:
  1. 1.
    We begin by initializing the model with the parameters obtained from the MDVRP pre-training phase, leaving the original deep neural network architecture intact;
  2. 2.
    Next, we discard the layers of the model associated with customer-specific features, since demands and time windows are not part of the TSP. Only the layers responsible for encoding node coordinates are retained, which correspond to the depot-embedding layers in the MDVRP model. We also keep the layers responsible for encoding the edge features, as spatial relationships between nodes remain essential in the TSP. For clarity, when we say that these layers are discarded, we mean that they are removed from the model architecture rather than being merely ignored;
  3. 3.
    We also modify the decoder to remove the dynamic features tied to VRP constraints (vehicle load, elapsed time, route length and the open-route indicator);
  4. 4.
    Finally, the solution decoding process follows TSP logic. At each step, the agent selects the next node to be visited by the salesman, and the first visited node is revisited only at the end of the trajectory. The masking mechanism ensures that nodes already visited cannot be selected again.
For the fine-tuning phase, we load the parameters from the pre-trained MDVRP model and train it for an additional 20 epochs on randomly generated instances, of each specific VRP variant. We use the same hyper-parameters as in the initial training phase. The dynamics of the pre-training and fine-tuning stages of TuneNSearch are illustrated in Fig. 2. We note that besides the main variants described in Section 3.3, TuneNSearch can be extended to any combination of VRP constraints. For example, it can handle combinations like the MDVRP with time windows, open routes and backhauls, allowing it to tackle more complex routing problems.
@@ 4.4 Local search algorithm
Despite the recent interest in using neural-based techniques to solve VRPs, there remains a limited integration between operations research and machine learning methods in the current literature. Bridging this gap offers an opportunity to combine the strengths of both fields, enabling the development of more powerful and efficient algorithms. To refine the solutions obtained by the neural-based model, we designed an efficient local search algorithm employed after inference, inspired by different existing methods. In this section, we describe each component of our local search.
Search operators and granular neighborhood: The core of our local search algorithm lies in applying a set of different operators within restricted neighborhoods of predefined size. Following Vidal (2022), we define the neighborhood size as 20, which allows for an efficient exploration of the granular neighborhood. A granular neighborhood is a restricted subset of the solution space that is defined based on proximity criteria. Rather than evaluating moves across the entire problem domain, the search is confined to carefully selected neighborhoods of limited size, where meaningful improvements are more likely to be found. Moves are restricted to node pairs $(a,b)$, where $b$ is one of the 20 closest nodes to $a$. Each move is evaluated within different neighborhoods, and any improvement in the cost function is immediately applied. The search procedure terminates when all applicable operators and moves have been applied without yielding any further improvement in the cost function. We adopt the same set of operators as Wouda et al. (2024), which were selected due to their ability to effectively explore the solution space while maintaining computational efficiency. The first three operators focus on node-level modifications, attempting moves between pairs of nodes whenever one node lies within the restricted neighborhood of the other. In contrast, the last two operators function at the route level, exchanging or repositioning customers between routes irrespective of their granular neighborhood. Each operator targets different aspects of the solution structure, ensuring a well-balanced and diverse set of moves to improve the solution quality. The chosen operators include:
@@
    (X, M)-exchange: involves swapping $X$ customers from a route (starting at a designated node), with a segment of $M$ customers (where $0\leq M\leq X$) starting at a different node. Importantly, the two segments must not overlap, ensuring that the exchange modifies the segments without duplicating any customers.
@@
    RELOCATE*: identifies and executes the best possible relocation of a single customer between two routes. It removes the customer from its current route and inserts it into the best position in the other route, such that the total cost is minimized (with respect to all other possible relocation moves).
@@
    SWAP*: evaluates and executes the most beneficial exchange of two customers between two routes, positioning each customer in the best possible location within the other route. Unlike a traditional swap, this operator exchanges two customers from different routes by inserting them into any position on the opposite route, rather than directly swapping them in place.
Local search algorithm workflow: We now describe the overall workflow of the proposed local search algorithm (see Algorithm 2). It begins with an initial solution $\tau$, which is the best solution obtained from the E-GAT model. The algorithm assumes a fixed number of iterations $I$, and solutions are represented as sequences of visited nodes. The best distance ($BD$) is initialized as the cost of $\tau$, and the best solution ($BS$) is set to $\tau$. Running an iterated local search starting from a randomly initialized solution can be very computationally expensive, and may require many iterations to converge. By seeding the search with a high-quality solution generated by the E-GAT model, we can significantly reduce the search space, and thus concentrate the search on more promising regions, accelerating convergence (see Section 5.4 for a comparison with randomly and greedily initialized solutions).
Before the iterative process starts, the algorithm explores the neighborhood of $\tau$ using the $Search$ function. This function applies all the operators and moves explained above to identify potential improvements to the solution. A more detailed explanation of this function is given below, in Algorithm 3. If a better solution is found during this step, both $BD$ and $BS$ are updated accordingly. After that, in each iteration, to escape local minimum, an offspring solution ($OS$) is generated using the $Crossover$ function. This function takes as parents the current $BS$ and a randomly generated solution, created with the $MakeRandom$ function. Here, we use the selective route exchange crossover (SREX) operator proposed by Nagata and Kobayashi (2010). Then, the neighborhood of $OS$ is explored through the $Search$ function. If the resulting solution is infeasible, the algorithm repairs it using the $Fix$ function, which applies the search operators again while considering only feasible moves. Finally, if the new solution’s cost is better than the current $BD$, the algorithm updates $BD$ and $BS$. This process is repeated for $I$ iterations.
The rationale behind the choice of always iterating the local search over $BS$ stems from balancing diversification and intensification. On the diversification side, the crossover operator combines the current best solution with a randomly generated solution, introducing more diversity and enabling the method to escape local optima. On the intensification side, anchoring the search around the best solution ensures that exploration remains focused on high-quality regions of the solution space, promoting steady progress toward better solutions. Iterating only from $BS$ risks getting trapped in local minimum, while iterating solely from random solutions would often lead the search into poor-quality regions. By combining both, we allow the algorithm to explore new structural patterns while preserving the advantages of high-quality routes already identified.
@@
[…]
@@
19:return BSBS
Search function: Algorithm 2 provides a high-level view of the entire local search procedure. A key component of this procedure is the $Search$ function, which explores different neighborhoods and applies the operators described earlier. Following the design of Vidal (2022), the implementation of $Search$ is detailed in Algorithm 3. The function takes a solution $\tau$ as input and iteratively improves it into a solution $\tau^{\prime}$. The search focuses first on node-level operators, i.e., (X, M)-exchange, MoveTwoClientsReversed and 2-OPT, which are applied in a randomized order at each function call to introduce stochasticity. For every node $a$ in the solution, the algorithm iterates over each node $b$ in its granular neighborhood, and evaluates each of the three operators on all admissible node pairs $(a,b)$. Whenever a move yields an improvement in the cost function, it is applied immediately and the solution is updated accordingly. This process repeats until a complete pass over all node pairs and operators produces no further improvements, at which point the node-level search terminates. Next, the function applies route-level operators (RELOCATE* and SWAP*), again in a different random order at each function call. It iterates over all route pairs in the solution and evaluates both operators. As before, any improving move is applied immediately and the solution is updated. Like the node-level search, this phase ends once all route pairs have been evaluated without improvement, completing the overall search process. Overall, the $Search$ function terminates only when no further locally improving move can be found.
Algorithm 3 Search Function
0: Solution $\tau$
0: Improved solution $\tau^{\prime}$
1: Initialize $\tau^{\prime}\leftarrow\tau$
2: Randomly shuffle the order of node and route operators
3:// Node-level search
4:repeat
5:  improvement $\leftarrow$ false
6:  for each customer node $a$ in $\tau^{\prime}$ do
7:   for each node $b$ in the granular neighborhood of $a$ do
8:    for each node operator nopnop do
9:     if $nop(a,b)$ improves cost then
10:      Apply $nop(a,b)$ to $\tau^{\prime}$ and update solution
11:      improvement $\leftarrow$ true
12:     end if
13:    end for
14:   end for
15:  end for
16:until improvement = false
17:// Route-level search
18:repeat
19:  improvement $\leftarrow$ false
20:  for each pair of routes $(\mathcal{R}_{i},\mathcal{R}_{j})$ in $\tau^{\prime}$ do
21:   for each route operator roprop do
22:    if $rop(\mathcal{R}_{i},\mathcal{R}_{j})$ improves cost then
23:     Apply $rop(\mathcal{R}_{i},\mathcal{R}_{j})$ to $\tau^{\prime}$
24:     Update solution and update solution
25:     improvement $\leftarrow$ true
26:    end if
27:   end for
28:  end for
29:until improvement = false
30:return $\tau^{\prime}$
Fix function: Within the local search algorithm workflow, the $Search$ function may occasionally produce infeasible solutions. This happens because the cost of a solution is evaluated using penalty coefficients for constraint violations. When constraints like vehicle capacity, time windows or route duration limits are violated, they are accounted for in the resulting cost function. However, in the case of the $Search$ function, we use very small penalty coefficients. As a result, infeasible solutions may be, on many occasions, more favorable than feasible ones, since the penalties are often outweighed by the reduction in the total distance traveled. Allowing these temporary infeasibilities enables the search process to explore beyond strictly feasible regions of the solution space, reaching areas that would otherwise remain inaccessible.
The $Fix$ function was introduced in our algorithm to repair these occasionally generated infeasible solutions. It does so by applying all previously described search operators to the infeasible solution, but this time with very large penalty coefficients (similar to prior studies, such as Vidal (2022)). By imposing overly high penalties for constraint violations, the $Fix$ function makes the cost of infeasible solutions highly uncompetitive, ensuring that only feasible moves are applied. Therefore, all results reported in this paper are guaranteed to be feasible. Furthermore, even after applying the $Fix$ function, additional feasibility checks are performed over the solution it produces to ensure that only feasible solutions can be returned by our local search algorithm.
Crossover: The crossover operator is central to our search, as it must balance two competitive objectives: exploring new regions of the search space, while also preserving promising solution structures. To achieve this balance, we use the SREX operator, whose design explicitly support both goals. This operator works as follows:
  1. 1.
    Two parents, $P_{A}$ and $P_{B}$, are given (in our case, the current best solution $BS$ and a randomly generated solution).
  2. 2.
    Two offspring solutions, $OS_{1}$ and $OS_{2}$, are initialized as copies of $P_{A}$.
  3. 3.
    Two sets of routes, $S_{A}$ (from $P_{A}$) and $S_{B}$ (from $P_{B}$), are randomly chosen.
  4. 4.
    The routes present in $S_{A}$ are removed from $OS_{1}$. Customers that appear in $S_{B}$ but not in $S_{A}$ are also removed from $OS_{1}$.
  5. 5.
    The routes present in $S_{A}$ are removed $OS_{2}$. The routes from $S_{B}$ are inserted into $OS_{1}$.
  6. 6.
    The routes from $S_{B}$ are inserted into $OS_{2}$, while excluding from $S_{B}$ any customers that are in $S_{B}$ but not in $S_{A}$.
  7. 7.
    Any remaining unserved customers are reinserted into each offspring based on the least costly insertions.
  8. 8.
    Evaluate both offsprings created and return the one with the better objective value.
For comparison, the order crossover (OX), used for example in many well-known studies (Prins, 2004; Vidal et al., 2013; İlhan İLHAN, 2021), creates an offspring by copying a single segment of nodes from one parent and filling the remaining positions in the order they appear in the other parent. Although conceptually simple and straightforward, OX often disrupts effective routes, resulting in offspring of limited quality. In contrast, SREX generates offsprings by inheriting multiple different routes that are kept mostly intact. This ensures that high-quality routes are more likely to be retained in the offspring, rather than being discarded or disrupted. As a result, SREX produces offsprings that remain close to favorable areas of the search space, while still introducing enough diversity to help the algorithm escape local minimum.
Beyond this theoretical advantage, SREX has also demonstrated strong empirical performance. Originally developed for the pickup and delivery problem with time windows, a task that closely resembles the VRP with backhauls and time windows, it demonstrated impressive performance compared to other well-known heuristics (Bent and Hentenryck, 2006; Røpke and Pisinger, 2006), improving 146 out of 298 benchmark problems. Its design also makes it highly flexible, allowing it to be applied across a wide range of VRP variants.
To further validate this choice, during the design phase of TuneNSearch, we conducted the experiments reported in Section 5.1, comparing the performance of our local search algorithm when using SREX versus OX. SREX consistently outperformed OX across all VRP variants, reinforcing its suitability in our work. For the aforementioned reasons, we rely on SREX in our algorithm, using OX solely for the TSP, as it is a problem in which solutions consist of a single sequence of nodes rather than distinct routes, making SREX inapplicable.
Fig. 3 depicts an overview of the complete inference process of TuneNSearch, coupled with the local search algorithm.
Figure 3: Overview of the inference process of TuneNSearch.